OpenStudy (anonymous):

Find the solution to the equation 64^(3 – x) = 4^2x. Using complete sentences, explain the procedure used to solve this equation.

OpenStudy (anonymous):

I suggest using logarithms

myininaya (myininaya):

I would write both sides with same base first since 4^3=64. 4^{3(3-x)}=4^{2x} Now since both bases are the same, I would equate the exponents, 3(3-x)=2x. you should be able to do the rest.